# On Secret-Message Transmission by Echoing Encrypted Probes

### v1

@@ III-G1 Comparison to $C_{key}$
$C_{key}=\log\left(1+\frac{S_{BA}}{S_{EA}+1}\right)=\log\left(1+\frac{a}{\alpha a+1}\right)$
$=\log\frac{A_{2}}{A_{1}}.$ (53)
It follows from (50) that
$\lim_{b\to\infty}C_{s,G}=\left(\log\left(1+\frac{1}{A_{1}}\right)-\log\left(1+\frac{1}{A_{2}}\right)\right)^{+}$
@@
Since $C_{s,G}$ increases with $b$ for $\beta>1$, then for $\beta>1$ we have $C_{s,G}<C_{key}$ for all $\alpha$, $a$ and $b$. This is expected as discussed before.
@@
$\lim_{a\to\infty}(\lim_{b\to\infty}C_{s,G})$
$=\lim_{a\to\infty}\left(\log\frac{A_{2}(A_{1}+1)}{A_{1}(A_{2}+1)}\right)^{+}$
@@
So, if both $a$ and $b$ are large but $b$ dominates $a$, then $C_{s,G}=C_{key}$. This is a special case of Proposition 2.
@@ IV-A Operations at Alice and Bob
In phase 2 of P-STEEP, Bob applies PSK nonlinear encryption, i.e., he sends out $\sqrt{p_{B}}x_{B}=\sqrt{p_{B}}e^{j\phi}r_{B}$ where $\phi$ is a secret phase value (meant for Alice) randomly chosen from the same discrete constellation as $\theta$. Here the construction of $x_{B}=e^{j\phi}r_{B}$ is different from that for G-STEEP with $n_{A}=n_{B}=1$. This nonlinear encryption fits naturally with PSK (a nonlinear modulation).
@@
Then Alice receives
@@
where $H(\phi)=\log M$ (the entropy of $\phi$). To determine $H(\phi|r_{A})$, we can view $\phi$ given $r_{A}$ as the optimal decision of $\phi$ from $r_{A}$.
@@
For $M=2^{m}$ with $m\geq 2$, the optimal decision of $\phi$ from $r_{A}$ takes approximately three possible values with the probabilities $1-p_{e,A}$, $\frac{1}{2}p_{e,A}$ and $\frac{1}{2}p_{e,A}$ respectively. In this case, we can write
$C_{A|B,P}=m-p_{e,A}-h_{2}(p_{e,A})$ (66)
@@ IV-B Operations at Eve
Since $\{r_{EA},r_{EB}\}$ is a one-to-one function of $\{r_{EA},r_{E}\}$, and $r_{EA}$ is independent of $r_{E}$ and $\phi$, we now know that $r_{E}$ is a sufficient statistic from $\{r_{EA},r_{EB}\}$ for $\phi$.
@@
Similar to $C_{A|B,P}$ in (65) and (66), we can express the effective capacity $C_{E|B,P}$ from Bob to Eve relative to $\phi$ as
@@
IV-C Secrecy capacity
The secrecy capacity of P-STEEP is
$C_{s,P}\doteq(C_{A|B,P}-C_{E|B,P})^{+}=h_{2}(p_{e,E})-h_{2}(p_{e,A})$ (76)
@@ IV-C Secrecy capacity
IV-D Ratio of bit error rates
For large $a$ and $b$, both $p_{e,A}$ and $p_{e,E}$ are small subject to $\alpha>1$ and $\beta>1$. In this case, $C_{s,P}$ is only a small positive value under (78). But the ratio $\gamma_{p}$ of $p_{e,A}$ over $p_{e,E}$ is also a meaningful metrics subject to a sufficiently long packet (e.g., a packet of $n$ independent bits with $np_{e,E}\approx 1$).
@@ V STEEP for Multiple Access (M-STEEP)
In phase 1 of M-STEEP, AP broadcasts a sequence of independent realizations of the random probing vector $\sqrt{p_{A}/n_{A}}\mathbf{x}\in\mathcal{C}^{n_{A}\times 1}$ with $\mathbf{x}$ being $\mathcal{CN}(\mathbf{0},\mathbf{I})$. Then UEi receives
@@
$=\sqrt{1/2}(\hat{p}_{i}+s_{i})\mathbf{h}_{Ai}^{\prime}+\mathbf{w}_{Ai}\in\mathcal{C}^{n_{A}\times 1}$ (86)
@@
It follows from (III-C) with $n_{B}=1$ or from (44) that the effective capacity from UEi to AP relative to $s_{i}$ is
$C_{A|i}\doteq I(s_{i};\hat{s}_{i})=\log\left(1+\frac{\frac{S_{Ai}}{2}}{\frac{S_{i}S_{Ai}/2}{(S_{i}+1)^{2}}+1}\right).$ (87)
where $\hat{s}_{i}$ is the MMSE estimate of $s_{i}$ by AP, $S_{Ai}=p_{ui}\|\mathbf{h}_{Ai}\|^{2}$ and $S_{i}$ was defined before.
V-A Effective Return Channel from each UE to Eve
@@
An achievable secrecy rate of M-STEEP from an arbitrarily selected UE1 to the AP relative to the message symbol $s_{1}$ from UE1 is
$C_{s,1}=(C_{A|1}-C_{E|1})^{+}=\left[\log\left(1+\frac{\frac{S_{A,1}}{2}}{\frac{S_{1}S_{A,1}/2}{(S_{1}+1)^{2}}+1}\right)\right.$
$\,\,\left.-\log\left(1+\frac{S_{E,1}/2}{(\gamma_{1}-1)S_{E,1}/2+1}\right)\right]^{+}$ (106)
@@
This proposition follows from (87) and (V-A). If $M=1$, it reduces to Corollary 1.
@@
Also $t_{1,M}=0$ for $M=1$, and $t_{1,M}$ for $M\geq 2$ is defined in (162) which is a function of $S_{E,A}$ and $S_{E,i}$ for all $i\neq 1$. And $t_{1,M}<\min(M-1,\alpha_{1}S_{1}+1)$. Consequently, $C_{s,1}>0$ if and only if
@@
This proposition has also been validated by computer simulations. If $M=1$, (108) reduces to (3). But more importantly, we see from (108) that for any given $M$, the secrecy rate from any UE to AP stays positive if that UE uses a sufficiently large power according to (108).
V-B Total Secrecy Capacity of M-STEEP
A total secrecy rate of M-STEEP can be expressed as
$C_{s}\doteq C_{s,1}+C_{s,2|1}+\cdots+C_{s,M|1,\cdots,M-1}$ (109)
with $C_{s,i|1,\cdots,i-1}\doteq[C_{A,i|1,\cdots,i-1}-C_{E,i|1,\cdots,i-1}]^{+}$ which is the secrecy rate from EUi to AP subject to $s_{1},\cdots,s_{i-1}$ being known to AP and Eve. Furthermore, we have $C_{A,i|1,\cdots,i-1}=I(s_{i};\mathbf{y}_{A}|s_{1},\cdots,s_{i-1},\mathbf{x})$ and $C_{E,i|1,\cdots,i-1}=I(s_{i};\mathbf{y}_{E}|s_{1},\cdots,s_{i-1})$. Here $\mathbf{y}_{A}=[\mathbf{y}_{A1}^{T},\cdots,\mathbf{y}_{AM}^{T}]^{T}$ and $\mathbf{y}_{E}=[\mathbf{y}_{E1}^{T},\cdots,\mathbf{y}_{EM}^{T},\mathbf{y}_{EA}^{T}]^{T}$.
It is easy to verify that the capacity from UEi to AP for $s_{i}$ is invariant to $s_{1},\cdots,s_{i-1}$ being known to AP or not. Hence $C_{A,i|1,\cdots,i-1}=C_{A,i}$ as given by (87).
The MSE of the MMSE estimate of $s_{i}$ by Eve using $\mathbf{y}_{E}$ and $s_{1},\cdots,s_{i-1}$ is
$\sigma_{\Delta s_{i,E|1,\cdots,i-1}}^{2}=1-\mathbf{r}_{i}^{H}\mathbf{R}_{i}^{-1}\mathbf{r}_{i}$ (110)
where $\mathbf{r}_{i}^{H}=\mathbb{E}\{s_{i}\mathbf{y}_{E|1,\cdots,i-1}^{H}\}$ and $\mathbf{R}_{i}=\mathbb{E}\{\mathbf{y}_{E|1,\cdots,i-1}\mathbf{y}_{E|1,\cdots,i-1}^{H}\}$. Here
$\mathbf{y}_{E|1,\cdots,i-1}\doteq$
@@ V-B Total Secrecy Capacity of M-STEEP
with $\mathbf{\bar{y}}_{E,j}=\mathbf{y}_{E,j}-\mathbf{h}_{Ej}^{\prime}s_{j}$. Furthermore, $\mathbf{r}_{i}$ is given by (91), and $\mathbf{R}_{i}$ is the same as $\mathbf{R}$ in (92) except that the $j$th diagonal block $\mathbf{R}_{j,j}$ of $\mathbf{R}$ for $j=1,\cdots,i-1$ should be replaced by $\mathbf{\bar{R}}_{j,j}=c_{j}\mathbf{h}_{Ej}^{\prime}\mathbf{h}_{Ej}^{\prime H}+\mathbf{I}_{n_{E}}$. Finally, $C_{E,i|1,\cdots,i-1}=-\log\sigma_{\Delta s_{i,E|1,\cdots,i-1}}^{2}$.
Subject to i.i.d. Gaussian random channel parameters, numerical results [30] suggest that the medium of $C_{s}$ increases as $M$ increases while the medium of $C_{s,1}$ decreases (but rather slowly) as $M$ increases. Further mathematical analysis of $C_{s}$ remains open.

### v2

@@ III-H2 Comparison to $C_{key}$
$C_{key}=\log\left(1+\frac{S_{BA}}{S_{EA}+1}\right)=\log\left(1+\frac{a}{\alpha a+1}\right)=\log\frac{A_{2}}{A_{1}}.$ (60)
Proposition 4
For any given $\alpha$ and $a$, there is a sufficiently large (but finite) $b$ and a sufficiently small (positive) $\beta$ such that $C_{key}<R_{s,G}$. But if $\beta\geq 1$, then $C_{key}>R_{s,G}$ for any finite $\alpha$, $a$ and $b$.
Proof:
Let $\gamma>1$ be such that $\gamma A_{2}-A_{1}<1$, i.e., $\gamma<\frac{1-A_{1}}{A_{2}}$. The first term of $R_{s,G}$ in (57) is strictly larger than $C_{key}+\log\gamma$ if $\frac{b}{2}>\frac{\gamma\frac{A_{2}}{A_{1}}-1}{1-(\gamma A_{2}-A_{1})}$. The second term of $R_{s,G}$ in (57) is smaller than $\log\gamma$ if $\beta<\frac{\gamma-1}{\frac{b}{2}(1-(\gamma-1)A_{2})}$. This proves the first statement in the proposition. To prove the second statement, first consider $\beta=1$. In this case, “$C_{key}>R_{s,G}$” is equivalent to
$\frac{A_{2}}{A_{1}}-1>\frac{\frac{b}{2}\left(1-\frac{A_{2}}{A_{1}}\right)}{\left(\frac{b}{2}A_{1}+1\right)\left(\frac{b}{2}A_{2}+1\right)}$ (61)
which always holds since the left side of (61) is positive and the right side of (61) is negative. Finally, notice that $R_{s,G}$ is a decreasing function of $\beta$. ∎
Alternatively, it follows from (57) that
$\lim_{b\to\infty}R_{s,G}=\left(\log\left(1+\frac{1}{A_{1}}\right)-\log\left(1+\frac{1}{A_{2}}\right)\right)^{+}$
@@
Since $R_{s,G}$ increases with $b$ for $\beta>1$, then for $\beta>1$ we have $R_{s,G}<C_{key}$ for all $\alpha$, $a$ and $b$, which is consistent with Proposition 4.
@@
$\lim_{a\to\infty}(\lim_{b\to\infty}R_{s,G})=\lim_{a\to\infty}\left(\log\frac{A_{2}(A_{1}+1)}{A_{1}(A_{2}+1)}\right)^{+}$
@@
In a practical term, we can say that if both $a$ and $b$ are large while $b$ dominates $a$, then $R_{s,G}\approx C_{key}$. This is a special case of Proposition 2.
@@ IV-A Description of P-STEEP
In phase 2 of P-STEEP, Bob applies PSK nonlinear encryption, i.e., he sends out $\sqrt{p_{B}}x_{B}=\sqrt{p_{B}}e^{j\phi}r_{B}$ where $\phi$ is a secret phase value (meant for Alice) randomly chosen (in this paper) from the same discrete constellation as $\theta$. Here the construction of $x_{B}=e^{j\phi}r_{B}$ is different from that for G-STEEP with $n_{A}=n_{B}=1$. This nonlinear encryption fits naturally with PSK (a nonlinear modulation).
@@
IV-B Analysis of the signal received by Alice in phase 2
The signal received by Alice in phase 2 is
@@ IV-B Analysis of the signal received by Alice in phase 2
Lemma 3
If $A$ is a circular complex Gaussian random variable with zero mean and variance $\sigma^{2}$, i.e., $\mathcal{CN}(0,\sigma^{2})$, then so is $e^{j\theta}A$ for any $\theta$. If $A$ and $B$ are two independent circular complex Gaussian random variables with zero means and variances $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$ respectively, then so are $e^{j\theta_{A}}A$ and $e^{j\theta_{B}}B$ for any $\theta_{A}$ and $\theta_{B}$.
Proof:
Since $A$ is $\mathcal{CN}(0,\sigma^{2})$, its amplitude $|A|$ and phase $\angle A$ are independent variables with $|A|$ being Rayleigh distributed and $\angle A$ being uniform distributed within $[0,2\pi)$. For any $\theta$, the modulo sum $(\theta+\angle A)_{mod-2\pi}$ remains uniform with $[0,2\pi)$. Hence the distributions of $|e^{j\theta}A|=|A|$ and $\angle(e^{j\theta}A)=(\theta+\angle A)_{mod-2\pi}$ do not change with $\theta$, i.e., $e^{j\theta}A$ is also $\mathcal{CN}(0,\sigma^{2})$. The other statement can be similarly proved (even if $\theta_{A}=\theta_{B}$). ∎
@@
where $H(\phi)=\log M$ (the entropy of $\phi$). To determine $H(\phi|r_{A})$, we can view $\phi$ given $r_{A}$ as the optimal decision (also known as hard decision) of $\phi$ from $r_{A}$.
@@
For $M=2^{m}$ with $m\geq 2$, the optimal decision of $\phi$ from $r_{A}$ at a high SNR (i.e., small $p_{e,A}$) takes approximately three possible values: the correct $\phi$ with the probability $1-p_{e,A}$, and the two nearest neighbors of $\phi$ with the probability $\frac{1}{2}p_{e,A}$ for each. In this case, we can write
$C_{A|B,P}\approx m-2\left(\frac{1}{2}p_{e,A}\log\frac{1}{\frac{1}{2}p_{e,A}}\right)$
$\,\,-(1-p_{e,A})\log\frac{1}{1-p_{e,A}}$
$=m-p_{e,A}-h_{2}(p_{e,A})$ (74)
@@ IV-C Analysis of the signals received by Eve in phases 1 and 2
Since $\{r_{EA},r_{EB}\}$ is a one-to-one function of $\{r_{EA},r_{E}\}$, and $r_{EA}$ is approximately independent of $r_{E}$ and $\phi$, we now know that $r_{E}$ is a sufficient statistic from $\{r_{EA},r_{EB}\}$ for $\phi$.
@@
Similar to $C_{A|B,P}$ in (73) and (IV-B), we can express the effective capacity $C_{E|B,P}$ from Bob to Eve relative to $\phi$ as
@@
IV-D Achievable secrecy rate
An achievable secrecy rate of P-STEEP is
$R_{s,P}\doteq(C_{A|B,P}-C_{E|B,P})^{+}=h_{2}(p_{e,E})-h_{2}(p_{e,A})$ (84)
@@ IV-D Achievable secrecy rate
IV-E Ratio of bit error rates
For large $a$ and $b$, both $p_{e,A}$ and $p_{e,E}$ are small subject to $\alpha>1$ and $\beta>1$. In this case, $R_{s,P}$ is only a small positive value under (86). But the ratio $\gamma_{p}$ of $p_{e,A}$ over $p_{e,E}$ is also a meaningful metrics subject to a sufficiently long packet (e.g., a packet of $n$ independent bits with $np_{e,E}\approx 1$).
@@ V STEEP for Multiple Access (M-STEEP)
V-A Description of M-STEEP
In phase 1 of M-STEEP, AP broadcasts a sequence of independent realizations of the random probing vector $\sqrt{p_{A}/n_{A}}\mathbf{x}\in\mathbb{C}^{n_{A}\times 1}$ with $\mathbf{x}$ being $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{A}})$. Then UEi receives
@@ V-A Description of M-STEEP
$=\sqrt{1/2}(\hat{p}_{i}+s_{i})\mathbf{h}_{Ai}^{\prime}+\mathbf{w}_{Ai}\in\mathbb{C}^{n_{A}\times 1}$ (94)
@@
V-B Effective channel from each UE to AP
It follows from (III-D) with $n_{B}=1$ or from (54) that an achievable rate from UEi to AP relative to $s_{i}$ (i.e., AP uses the signal from UEi and the original probing vector to extract the information from $s_{i}$) is
$R_{A|i}\doteq I(s_{i};\hat{s}_{i})=\log\left(1+\frac{\frac{S_{Ai}}{2}}{\frac{S_{i}S_{Ai}/2}{(S_{i}+1)^{2}}+1}\right),$ (95)
where $\hat{s}_{i}$ is the MMSE estimate of $s_{i}$ by AP from $\mathbf{y}_{Ai}$ and $\mathbf{x}$, $S_{Ai}=p_{ui}\|\mathbf{h}_{Ai}\|^{2}$ and $S_{i}$ was defined before.
Note that when $M\geq 2$, $R_{A|i}$ is a lower bound on $C_{A|i}\doteq I(s_{i};\mathbf{y}_{A}|\mathbf{x})$ with $\mathbf{y}_{A}=[\mathbf{y}_{A1}^{T},\cdots,\mathbf{y}_{AM}^{T}]^{T}$, the latter of which represents the optimal effective channel capacity from UEi to AP. However, even with $R_{A|i}$, we still can show that M-STEEP achieves a positive secrecy rate for each UE.
V-C Effective channel from each UE to Eve
@@ V-D Achievable secrecy rate from each UE to AP
An achievable secrecy rate of M-STEEP from an arbitrarily selected UE1 to the AP relative to the message symbol $s_{1}$ from UE1 is $R_{s,1}=(\tilde{R}_{s,1})^{+}$ with
$\tilde{R}_{s,1}\doteq I(s_{1};\mathbf{y}_{A}|\mathbf{x})-I(s_{1};\mathbf{y}_{E})\geq I(s_{1};\mathbf{y}_{A1}|\mathbf{x})-I(s_{1};\mathbf{y}_{E})$
$=R_{A|1}-C_{E|1}=\log\left(1+\frac{\frac{S_{A,1}}{2}}{\frac{S_{1}S_{A,1}/2}{(S_{1}+1)^{2}}+1}\right)$
$\,\,-\log\left(1+\frac{S_{E,1}/2}{(\gamma_{1}-1)S_{E,1}/2+1}\right),$ (114)
@@
This proposition follows from (95) and (V-C). If $M=1$, it reduces to Corollary 1.
@@
Also $t_{1,M}=0$ for $M=1$, and $t_{1,M}$ for $M\geq 2$ is defined in (179) which is a function of $S_{E,A}$ and $S_{E,i}$ for all $i\neq 1$. And $t_{1,M}<\min(M-1,\alpha_{1}S_{1}+1)$. Consequently, $R_{s,1}>0$ if and only if
@@
This proposition has also been validated by computer simulations. If $M=1$, (116) reduces to (58). But more importantly, we see from (116) that for any given $M$, the secrecy rate from any UE to AP stays positive if that UE uses a sufficiently large power according to (116). Specifically, $1-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}$ in (116) is virtually invariant to a moderate $M$ (e.g., in order of tens) when $\alpha_{1}S_{1}$ is large (e.g., in order of 30dB).
V-E Total (or sum) secrecy rate of M-STEEP
V-E1 A general expression
Now define $\mathbf{s}=[s_{1},\cdots,s_{M}]^{T}$ and recall $\mathbf{y}_{A}=[\mathbf{y}_{A1}^{T},\cdots,\mathbf{y}_{AM}^{T}]^{T}$ and $\mathbf{y}_{E}=[\mathbf{y}_{E1}^{T},\cdots,\mathbf{y}_{EM}^{T},\mathbf{y}_{EA}^{T}]^{T}$. A total achievable secrecy rate of M-STEEP from all UEs to AP (in bits per $M+1$ complex channel uses) can be written as $\tilde{R}_{s}^{+}$ with
$\tilde{R}_{s}=I(\mathbf{s};\mathbf{y}_{A}|\mathbf{x})-I(\mathbf{s};\mathbf{y}_{E}),$ (117)
where the condition on $\mathbf{x}$ in the first term is because of $\mathbf{x}$ being known to AP. This expression (117) is a straightforward extension of the theory behind (6) or (II-B). In this case, $\mathbf{K}_{s}=\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}$, and the secret message from each UE is chosen independently (or locally) while the WTC coding scheme can be centrally designed and publicly shared.
Since $I(\mathbf{s};\mathbf{y}_{A}|\mathbf{x})=\sum_{i=1}^{M}\tilde{R}_{s,i|A}$ with $\tilde{R}_{s,i|A}=I(s_{i};\mathbf{y}_{A}|s_{1},\cdots,s_{i-1},\mathbf{x})$, and $I(\mathbf{s};\mathbf{y}_{E})=\sum_{i=1}^{M}\tilde{R}_{s,i|E}$ with $\tilde{R}_{s,i|E}=I(s_{i};\mathbf{y}_{E}|s_{1},\cdots,s_{i-1})$, we can write
$\tilde{R}_{s}=\sum_{i=1}^{M}\tilde{R}_{s,i}$ (118)
with $\tilde{R}_{s,i}=\tilde{R}_{s,i|A}-\tilde{R}_{s,i|E}$. Here $\tilde{R}_{s,i}$ could be negative and its value in general depends on the ordering of the UEs. But $\tilde{R}_{s}$ is invariant to the ordering. Also note that $\tilde{R}_{s,1}$ here is the same as (5).
Furthermore, we know
$\tilde{R}_{s,i|A}\geq I(s_{i};\mathbf{y}_{Ai}|s_{1},\cdots,s_{i-1},\mathbf{x})$
$=I(s_{i};\mathbf{y}_{Ai}|\mathbf{x})=R_{A|i},$ (119)
which is given by (95). And
$\tilde{R}_{s,i|E}=-\log\sigma_{\Delta s_{i,E|1:i-1}}^{2}$ (120)
where $\sigma_{\Delta s_{i,E|1:i-1}}^{2}$ is the MSE of the MMSE estimate of $s_{i}$ by Eve using $\mathbf{y}_{E}$ and $s_{1},\cdots,s_{i-1}$ (as if they are known to Eve), i.e.,
$\sigma_{\Delta s_{i,E|1:i-1}}^{2}=1-\mathbf{r}_{i}^{H}\mathbf{R}_{i}^{-1}\mathbf{r}_{i}$ (121)
where $\mathbf{r}_{i}^{H}=\mathbb{E}\{s_{i}\mathbf{y}_{E|1:i-1}^{H}\}$ and $\mathbf{R}_{i}=\mathbb{E}\{\mathbf{y}_{E|1:i-1}\mathbf{y}_{E|1:i-1}^{H}\}$. Here
$\mathbf{y}_{E|1:i-1}\doteq$
@@ V-E1 A general expression
with $\mathbf{\bar{y}}_{E,l}=\mathbf{y}_{E,l}-\mathbf{h}_{El}^{\prime}s_{l}$ and $l=1,\cdots,i-1$. Furthermore, $\mathbf{r}_{i}$ is given by (99), and $\mathbf{R}_{i}$ is the same as $\mathbf{R}$ in (100) except that the $l$th diagonal block $\mathbf{R}_{l,l}$ of $\mathbf{R}$ for $l=1,\cdots,i-1$ should be replaced by $\mathbf{\bar{R}}_{l,l}=\frac{1}{2}c_{l}\mathbf{h}_{El}^{\prime}\mathbf{h}_{El}^{\prime H}+\mathbf{I}_{n_{E}}$.
Next we show a case where all terms in (118) can be made positive by a power control even when Eve’s receive channels from all other nodes are stronger than those among AP and UEs.
V-E2 A special case
To gain further insights into $\tilde{R}_{s}$, let us assume $n_{A}=n_{B}=n_{E}=1$. All channel gains can be now lumped into noise variances. Specifically, in phase 1 of M-STEEP, the AP broadcasts the probe $p$, UEi receives $y_{i}=p+w_{i}$ for $i=1,\cdots,M$, and Eve receives $y_{E,A}=p+w_{E,A}$. Let $\hat{p}_{i}$ be the MMSE estimate of $p$ by UEi from $y_{i}$. In phase 2 of M-STEEP, UEi sends $x_{i}=\hat{p}_{i}+s_{i}$ to the AP using the $i$th orthogonal channel, and hence the AP receives $y_{A,i}=x_{i}+w_{A,i}$ and Eve receives $y_{E,i}=x_{i}+w_{E,i}$ for all $i$. Assume $p$ and $s_{i}$ for all $i$ are i.i.d. $\mathcal{CN}(0,1)$ and all noises ($w_{i}$, $w_{A,i}$, $w_{E,A}$ and $w_{E,i}$) are i.i.d. $\mathcal{CN}(0,\sigma_{*}^{2})$ with * chosen according to the index of the noise.
It follows that $\hat{p}_{i}=(1-\mu_{i})y_{i}$ with $\mu_{i}=\frac{\sigma_{i}^{2}}{1+\sigma_{i}^{2}}$, and the MSE of the MMSE estimate of $s_{i}$ by the AP from $y_{A,i}$ and $p$ is
$\sigma_{\Delta s_{i}}^{2}=\frac{\mu_{i}(1-\mu_{i})+\sigma_{A,i}^{2}}{1+\mu_{i}(1-\mu_{i})+\sigma_{A,i}^{2}}.$ (123)
Hence $\tilde{R}_{s,i|A}\geq-\log\sigma_{\Delta s_{i}}^{2}$, the right-side of which is a special case of $R_{A|i}$ in (95) with $p_{ui}=\frac{2}{\sigma_{A,i}^{2}}$ and $\|\mathbf{h}_{Ai}\|=1$.
One can also verify that the MSE of the MMSE estimate of $s_{i}$ by Eve from $\mathbf{y}_{E}\doteq[y_{E,A},y_{E,1},\cdots,y_{E,M}]^{T}$ is
$\sigma_{\Delta s_{i|E}}^{2}=1-\mathbf{e}_{i}^{T}\left(\mathbf{A}-(1/b)\mathbf{c}\mathbf{c}^{T}\right)^{-1}\mathbf{e}_{i}$ (124)
where $\mathbf{e}_{i}$ consists of all zeros except for its $i$th element equal to one, $\mathbf{c}^{T}=[\mu_{1}^{\prime},\cdots,\mu_{M}^{\prime}]$, $\mu_{i}^{\prime}=1-\mu_{i}$, $b=1+\sigma_{E,A}^{2}$, and
$\mathbf{A}=\left[\begin{array}[]{cccc}a_{E,1}&\mu_{1}^{\prime}\mu_{2}^{\prime}&\cdots&\mu_{1}^{\prime}\mu_{M}^{\prime}\\ \mu_{2}^{\prime}\mu_{1}^{\prime}&a_{E,2}&\cdots&\cdots\\ \cdots&\cdots&\ddots&\mu_{M-1}^{\prime}\mu_{M}^{\prime}\\ \mu_{M}^{\prime}\mu_{1}^{\prime}&\cdots&\mu_{M}^{\prime}\mu_{M-1}^{\prime}&a_{E,M}\end{array}\right]$ (125)
with $a_{E,i}=1+\mu_{i}^{\prime}+\sigma_{E,i}^{2}$. Furthermore, the MSE of the MMSE estimate of $s_{i}$ by Eve from $\mathbf{y}_{E}$ and $s_{1},\cdots,s_{i-1}$ is
$\sigma_{\Delta s_{i|E,1:i-1}}^{2}=1-\mathbf{e}_{i}^{T}(\mathbf{A}_{i}-(1/b)\mathbf{c}\mathbf{c}^{T})^{-1}\mathbf{e}_{i}$ (126)
where $\mathbf{A}_{i}$ is the same as $\mathbf{A}$ after its $l$th diagonal element being replaced by $\mu_{l}^{\prime}+\sigma_{E,l}^{2}$ for $l=1,\cdots,i-1$. Hence $\tilde{R}_{s,i|E}=-\log\sigma_{\Delta s_{i|E,1:i-1}}^{2}$.
Proposition 7
For a symmetric network where $\sigma_{i}=\sigma$, $\sigma_{A,i}=\sigma_{A}$ and $\sigma_{E,i}=\sigma_{E}$ for all $i=1,\cdots,M$, the total achievable secrecy rate in (117) is
$\tilde{R}_{s}=\tilde{R}_{s,1}+\tilde{R}_{s,2}+\cdots+\tilde{R}_{s,M}$
$\geq\log\frac{\sigma_{\Delta s_{1|E}}^{2}}{\sigma_{\Delta s_{1}}^{2}}+\log\frac{\sigma_{\Delta s_{2|E,1}}^{2}}{\sigma_{\Delta s_{2}}^{2}}+\cdots+\log\frac{\sigma_{\Delta s_{M|E,1:M-1}}^{2}}{\sigma_{\Delta s_{M}}^{2}}.$ (127)
Referring to the $i$th term after $\geq$” in (7) as $R_{s,i}^{\prime}$, we have $\tilde{R}_{s,i}\geq R_{s,i}^{\prime}$ for all $i$, and
$R_{s,1}^{\prime}>R_{s,2}^{\prime}>\cdots>R_{s,M}^{\prime}.$ (128)
If $\beta_{0}\doteq\frac{\sigma_{A}^{2}}{\sigma_{E}^{2}}\leq 1$ (i.e., Eve’s channels from UEs are not stronger than AP’s channels from UEs), $R_{s,M}^{\prime}>0$. If $\beta_{0}$ is fixed and larger than one (i.e., the opposite case from the above and virtually regardless of the channels in phase 1), and $\sigma_{A}^{2}$ is inversely proportional to the power $p_{B}$ used by each UE, then there is a finite threshold $\bar{p}_{B}$ such that $R_{s,M}^{\prime}>0$ if $p_{B}>\bar{p}_{B}$. For $\beta_{0}>1$ and a large $M$, $\bar{p}_{B}$ increases linearly with $M$.
Proof:
See Appendix -D. ∎
Note that $\tilde{R}_{s}$ in (118) corresponds to a scaled form of all-user (“individual or collective”) secrecy in [34], $\tilde{R}_{s,M}$ in (118) a scaled form of single-user “individual secrecy” in [34], and $\tilde{R}_{s,1}$ in (5) a scaled form of single-user “collective secrecy” in [34]. While the scheme considered in [34] and [35] subject to orthogonal access is not able to make $\tilde{R}_{s}>0$ when Eve’s channels are somewhat stronger than users, M-STEEP can make all components of $\tilde{R}_{s}$ (even the smallest $\tilde{R}_{s,M}$) positive by choosing strong enough powers from users in phase 2. However, a more complete study of the secrecy rate region subject to power constraints on both AP and users remains open. The assumptions in [34] and [35] do not cover such a round-trip collaboration exploited by STEEP.